OpenStudy (anonymous):

x^2-6x=8 Solve for x?

OpenStudy (anonymous):

I think I'm getting tired :(

OpenStudy (kinggeorge):

\[x^2 -6x-8 = 0\] and then use the quadratic formula

OpenStudy (anonymous):

Got it, thanks. Some of these processes are blending together. Thank you

OpenStudy (kinggeorge):

if you work through everything you get \[x = 3 \pm \sqrt17\]

OpenStudy (anonymous):

So next one is: \[\sqrt{x-6} = 3\]

OpenStudy (anonymous):

Same thing? Get in standard form first/

OpenStudy (kinggeorge):

exactly. Square both sides, and use the same strategy. In this case however you get \[\sqrt{x-6}=3 \rightarrow x-6=9 \rightarrow x=15\] so its more simple than the last.

OpenStudy (anonymous):

Wow...easier than it , thanks again